OpenStudy (isninajuma):

A white-tailed deer can sprint at speeds up to 30 miles per hour. American bison can run at speeds up to 3,520 feet per minute. Which animal is faster and by how many miles per hour? There are 5,280 feet in one mile.

OpenStudy (tkhunny):

Convert: \(\dfrac{3520\cdot ft}{1\;min}\cdot\dfrac{1\;mi}{5280\; ft}\cdot\dfrac{60\;min}{1\;hr}\) See how each new fraction is just one (1) and doesn't change the magnitude? We're just changing the units.